Lesiba invests an amount of money in an a ount earning 13,88% interest per year, ompounded weekly. After ve years, this amount has a umulated to R50 000. The amount that was invested initially is

[1] R26 105,54.

[2] R15 300,00.

[3] R29 515,94.

[4] R25 001,79.

[5] R34 700,00.


1
Expert's answer
2021-06-15T09:17:31-0400

P=principal or amount invested initially,t= time in years

A = amount , r = the interest rate per time period, and n = the number of time periods

t=5×52=260

"P=\frac{A}{(1+\frac{r}{n})^{t×n}}"

The amount that was invested initially is: 

"P = 50,000\/(1 + 0.1388\/52)^{260} = R25,001.79."

So, the correct answer is 4.